# FiniteTimeBibliography.bib

### v1

@@ 3 Quantum eigenspace thermalization with degeneracies
Now, we turn to the problem of quantum thermalization in a $D$-dimensional Hilbert space $\mathcal{H}$. As has long been recognized [vonNeumannThermalization, DAlessio2016, deutsch2018eth], it is impossible to have every conceivable observable thermalize (even o.a.) under Hamiltonian quantum dynamics with a complete orthonormal basis of energy levels $\lvert E_{n}\rangle$ (which we assume are indexed in ascending order, $E_{n+1}\geq E_{n}$). This is partly because a Hamiltonian quantum system is intrinsically non-ergodic, in the classical sense, in the Hilbert space [vonNeumannThermalization] due to conserving the overlaps $\lvert\langle E_{n}|\psi(t)\rangle\rvert^{2}$ of a state $\lvert\psi(t)\rangle$ with the energy eigenstates $\lvert E_{n}\rangle$.
@@
We will review two kinds of thermalization criteria for the observable for infinite time intervals, while setting up notation for the subsequent sections. One is the well known notion of eigenstate thermalization in the sense of Eq. (2), which implies thermalization o.a. given nondegenerate spectra (Sec. 3.1). Another criterion — which appears to be widely unknown (by our accounting, it has been discussed at least thrice before [Sunada, VenutiLiu, TumulkaEsT], but is not usually invoked in discussions of eigenstate thermalization in which nondegeneracy still plays a key role [DAlessio2016, MoriETHreview, deutsch2018eth, KawamotoStrongETH]) — is that in systems with degenerate spectra, one can impose thermalization in degenerate eigenspaces, from which thermalization in other states follows (Sec. 3.2). Our version is somewhat stricter than Refs. [VenutiLiu, TumulkaEsT], which will prove useful later. We discuss this phenomenon under the separate name of eigenspace thermalization to emphasize a conceptual distinction: in degenerate systems, an observable could satisfy eigenstate thermalization in a specific eigenbasis without eigenspace thermalization, and thereby fail to thermalize in most states. After discussing these notions, we will use eigenspace thermalization to qualitatively motivate energy-band thermalization in Sec. 3.3 (postponing a quantitative treatment to later sections), by arguing that energy levels within an energy band of resolution $\Delta E$ may be regarded as approximately degenerate.
3.1 Eigenstate thermalization $+$ non-degeneracy $\implies$ idealized thermalization
@@
To discuss eigenstate thermalization as a criterion for thermalization at a more technical level [vonNeumannThermalization, JensenShankarETH, deutsch1991eth, Deutsch2025Supplement, srednicki1994eth, srednicki1999eth, rigol2008eth, DAlessio2016, MoriETHreview, deutsch2018eth], we will find it convenient to restrict ourselves to a subspace $\Sigma_{d}\subseteq\mathcal{H}$ of $d$ energy levels to which initial states are assumed to belong, that we will formally call an energy shell (physically, $\Sigma_{d}$ may represent a narrow range of energies $E_{n}\in[E_{\min},E_{\max}]$, and accounts for initial states with support in this range [^9] For thermalization in a more general class of initial states with a narrow spread but with asymptotic tails outside a finite range, see [srednicki1999eth]; for the behavior of states without a narrow spread, see also [GenericETH]. We do not consider these more realistic cases for simplicity, and expect that the extension to such cases is straightforward along the lines of Ref. [srednicki1999eth], by expanding around energy shells.; it may also exclude some energy levels if desired, for example by restricting to an eigenspace of a conserved charge $\hat{Q}$ within this range). For energy shells of experimentally accessible energy widths $\Delta E$, we expect that $d\sim D/c_{\Sigma}$ where $c_{\Sigma}$ is some “accessible” constant [i.e., in the range of $Q$ in Eq. (4); see also Eq. (169)]. Thus, $d$ is usually an inaccessibly large quantity, comparable to $D$.
@@
Proposition 3.1 (Quantum eigenstate thermalization implies thermalization o.a. given a nondegenerate spectrum [vonNeumannThermalization, srednicki1999eth]).
@@
Though well known [vonNeumannThermalization, srednicki1999eth, DAlessio2016, MoriETHreview, deutsch2018eth], we briefly review the proof in App. D.2.1, as we will use similar arguments on other occasions in this paper. It is based on the following relation: If the energy levels $E_{n}$ within $\Sigma_{d}$ are nondegenerate, then for any observable $\hat{\Pi}_{A}:\mathcal{H}\to\mathcal{H}$ in the full system, and any density operator $\hat{\rho}:\Sigma_{d}\to\Sigma_{d}$ in the energy shell,
@@ 3.2 Eigenspace thermalization $\implies$ idealized thermalization
Figure 4: Schematic depiction of eigenspace thermalization in terms of the relevant matrix elements between pairs $(n,m)$ of the energy levels $(E_{n},E_{m})$; essentially, only the $E_{n}=E_{m}$ pairs are relevant. This matrix element structure can appear different between a spectrum with rare/small degeneracies (4(a)) and with frequent/large degeneracies (4(b)), and therefore more complicated than eigenstate thermalization. However, in terms of energy differences it corresponds to the extremely natural $\Delta E=0$ independent of the spectral details, and can imply thermalization without any knowledge of the spectrum. Even in this case, pairs with $\Delta E=0$ is still not accessible, and as we will argue in Sec. 4.2, eigenspace thermalization is not necessary for accessible thermalization especially in highly degenerate systems.
@@
for some small $\epsilon$. Such a relation has been used to define thermalization in eigenspaces in, e.g., Ref. [VenutiLiu].
While the $\lvert\psi\rangle$ are dynamically independent, they are not all linearly independent; it is therefore desirable to re-express Eq. (45) in terms of a simpler, more tractable criterion in $\mathcal{H}(E)$. For this purpose, we note that the Cauchy-Schwarz inequality (applied to the trace inner product $\operatorname{Tr}[\hat{P}^{\dagger}\hat{Q}]\leq\sqrt{\operatorname{Tr}[\hat{P}^{\dagger}\hat{P}]\operatorname{Tr}[\hat{Q}^{\dagger}\hat{Q}]}$ for $\hat{P}=\lvert\psi\rangle\langle\psi\rvert$ and $\hat{Q}=\hat{\Pi}(E)\left(\hat{\Pi}_{A}-\langle\hat{\Pi}_{A}\rangle_{\Sigma_{d}}\hat{\mathds{1}}\right)\hat{\Pi}(E)$) implies:
@@ 4.1 Weighted time averages
We will also take $w(t)\geq 0$. Among other things, through a relation between non-negativity conditions and Fourier transforms [FTpositivityConvex, FTpositivity2014], this implies that
@@
Finally, we need to formalize the notion of when $w(t)$ represents a “long-time average”, relative to the timescale set by the bandwidth $\Delta E$ For this, we want $w(t)$ to have appreciable magnitude at least over some long time interval $T\sim 2\pi/\Delta E$, such that $\widetilde{w}(\delta E)$ is appreciable only within $\lvert\delta E\rvert<\Delta E$. Specifically, for some small cutoff $10<w_{0}<$ 1, we require that $\widetilde{w}(\delta E)$ satisfies:
@@ 4.2 Thermalization in energy bands
$[\hat{A}]_{\Sigma_{d},\Delta E}\equiv\sum_{\begin{subarray}{c}n,m:\\ \lvert E_{n}\rangle,\lvert E_{m}\rangle\in\Sigma_{d}\\ \lvert E_{n}-E_{m}\rvert<\Delta E\end{subarray}}\lvert E_{n}\rangle\langle E_{n}\rvert\hat{A}\lvert E_{m}\rangle\langle E_{m}\rvert.$ (61)
@@
We can now define “energy-band thermalization” [^12] We note that this structure also occurs if one combines semiclassical results [ZelditchTransition, Sunada] on diagonal and off-diagonal matrix elements connected to classically ergodic systems, and formally removes the $\Delta E\to 0$ limit. by analogy with eigenspace thermalization [Definition 3.2], with the pair of eigenspace projectors $\hat{\Pi}(E)$ replaced by the energy-band selector in Eq. (59):
@@
However, for this to be nontrivial with “accessible” values of $\epsilon$, the number of distinct energy levels in the energy shell should be close to its dimension, $n\gg\epsilon^{2}d/\lambda^{2}$, so that $n_{\lambda}\ll n$ by the above constraint. In that case, we can claim “weak eigenspace thermalization” ($n_{\lambda}/n\ll 1$), in analogy with the notion of “weak ETH” [BiroliWeakETH, MoriWeakETH, MoriETHreview] where almost all (but not necessarily all) eigenstates thermalize; more discussion on the physical relevance of such “weak” notions will follow in the Conclusion, Sec. 9.
However, there is a complication: the number of eigenspaces is guaranteed to be accessible in an experiment only if $n$ itself is accessibly small [^13] This is because in a system with $n$ eigenspaces, the spectral form factor [Haake] (see also Sec. 8), used to measure energy level statistics [SFFmeas, pSFF], may oscillate at late times around values as small as $1/n$., such as $n<cN^{\alpha}$ for an $N$-particle system, which requires an unusually high level of degeneracy (close to $\Theta(d)$) in most energy levels. It is therefore difficult to establish that a system has the requisite high number of eigenspaces in most cases.
@@
    If one wants to keep to “accessible” values of $\epsilon$, one can only show weak eigenspace thermalization, that too in a (formally) restricted class of systems where the number of levels still scales as nearly $d$ (i.e. where a high degree of degeneracy is not typical), though this may cover several cases of interest (e.g., such as systems with a time reversal symmetry showing Kramers’ twofold degeneracy [Haake], but otherwise assumed to be nondegenerate).
@@ 4.3 Energy-band thermalization $\implies$ almost all physical states thermalize
for energy-band thermalization to guarantee thermalization with accuracy $\lambda$, provided that $\mu(\rho)$ is “accessible”. This is in fact anticipated by classical mechanics: such states are precisely the quantum analogue of the “bulky” states in Sec. 2. This analogy is quite strong [^14] We have sometimes been tempted to refer to this (old and well established) semiclassical relation as “Entanglement = Phase space volume”, with the qualifier that entanglement here means the exponential of the second Rényi entropy, following a set of conjectures in quantum gravity [SusskindCV] relating quantum information measures to more geometric (semi-)classical ones.: if a classical limit exists, the parameter $\mu(\rho)$ can be directly identified with the phase space volume of a classical region in which (say) $\rho$ has uniform support (if it is associated with such a classical state) [StechelMeasure, StechelHeller].
@@
=$\mathcal{B}=\{\lvert k\rangle\}_{k=0}^{d-1}\;\;\ \subset\Sigma_{d}.$ (71)
@@
=$\mathcal{B}=\left\{\lvert s_{1},\ldots,s_{N}\rangle:s_{j}\in\{{0,1}\}\right\},$ (72)
@@
Then, for any complete orthonormal basis of pure states $\mathcal{B}$ within the energy shell $\Sigma_{d}$, the fraction $f_{\lambda}[\mathcal{B}]$ of states that fail to thermalize o.a. to accuracy $\lambda$, i.e. [with $\overline{\Theta}(x)=1$ if $x$ is true and 00 otherwise]
@@
As long as $\epsilon,w_{0}\ll\lambda$, the fraction of basis states with non-thermal time averages to accuracy $\lambda$ (weighted by a given $w(t)$) is vanishingly small. But we cannot force $f_{\lambda}[\mathcal{B}]<1/d$ (which would ensure that all states thermalize) with accessible values of $\epsilon$ and $w_{0}$: there appears to be no accessible way to rule out a small fraction of non-thermal physical states in any accessible time average. These are a quantum analogue of measure-zero exceptions, such as periodic orbits in a classical system showing measure-theoretic ergodicity. As noted in Eq. (68), it is possible to rule out these features with “inaccessible” values of $\epsilon$ and $w_{0}$ by directly considering the thermalization of pure states given energy-band thermalization.
@@ 5 Global thermalization from a single initial state
To connect autocorrelators to global energy-band thermalization, we only need a straightforward adaptation of the proofs of wavefunction ergodicity [ZelditchTransition, Sunada, Zelditch] from global eigenstate thermalization to energy-band thermalization with a careful accounting of weighted time averages. This is carried out in Sec. 5.2. More significantly, in Sec. 5.3, we combine this result with those of Sec. 4 to completely bypass properties of the energy levels, and directly connect the decay of autocorrelators to global thermalization, realizing a fully quantum analogue of the classical statement containing Eq. (1).

### v2

@@ 3 Quantum eigenspace thermalization with degeneracies
Now, we turn to the problem of quantum thermalization in a $D$-dimensional Hilbert space $\mathcal{H}$. As has long been recognized [10, 17, 19], it is impossible to have every conceivable observable thermalize (even o.a.) under Hamiltonian quantum dynamics with a complete orthonormal basis of energy levels $\lvert E_{n}\rangle$ (which we assume are indexed in ascending order, $E_{n+1}\geq E_{n}$). This is partly because a Hamiltonian quantum system is intrinsically non-ergodic, in the classical sense, in the Hilbert space [10] due to conserving the overlaps $\lvert\langle E_{n}|\psi(t)\rangle\rvert^{2}$ of a state $\lvert\psi(t)\rangle$ with the energy eigenstates $\lvert E_{n}\rangle$.
@@
We will review two kinds of thermalization criteria for the observable for infinite time intervals, while setting up notation for the subsequent sections. One is the well known notion of eigenstate thermalization in the sense of Eq. (2), which implies thermalization o.a. given nondegenerate spectra (Sec. 3.1). Another criterion — which appears to be widely unknown (by our accounting, it has been discussed at least thrice before [41, 50, 66], but is not usually invoked in discussions of eigenstate thermalization in which nondegeneracy still plays a key role [17, 18, 19, 51]) — is that in systems with degenerate spectra, one can impose thermalization in degenerate eigenspaces, from which thermalization in other states follows (Sec. 3.2). Our version is somewhat stricter than Refs. [50, 66], which will prove useful later. We discuss this phenomenon under the separate name of eigenspace thermalization to emphasize a conceptual distinction: in degenerate systems, an observable could satisfy eigenstate thermalization in a specific eigenbasis without eigenspace thermalization, and thereby fail to thermalize in most states. After discussing these notions, we will use eigenspace thermalization to qualitatively motivate energy-band thermalization in Sec. 3.3 (postponing a quantitative treatment to later sections), by arguing that energy levels within an energy band of resolution $\Delta E$ may be regarded as approximately degenerate.
3.1 Eigenstate thermalization ++ non-degeneracy $\implies$ idealized thermalization
@@
To discuss eigenstate thermalization as a criterion for thermalization at a more technical level [10, 11, 12, 13, 14, 15, 16, 17, 18, 19], we will find it convenient to restrict ourselves to a subspace $\Sigma_{d}\subseteq\mathcal{H}$ of $d$ energy levels to which initial states are assumed to belong, that we will formally call an energy shell (physically, $\Sigma_{d}$ may represent a narrow range of energies $E_{n}\in[E_{\min},E_{\max}]$, and accounts for initial states with support in this range [^9] For thermalization in a more general class of initial states with a narrow spread but with asymptotic tails outside a finite range, see [15]; for the behavior of states without a narrow spread, see also [70]. We do not consider these more realistic cases for simplicity, and expect that the extension to such cases is straightforward along the lines of Ref. [15], by expanding around energy shells.; it may also exclude some energy levels if desired, for example by restricting to an eigenspace of a conserved charge $\hat{Q}$ within this range). For energy shells of experimentally accessible energy widths $\Delta E$, we expect that $d\sim D/c_{\Sigma}$ where $c_{\Sigma}$ is some “accessible” constant [i.e., in the range of $Q$ in Eq. (4); see also Eq. (169)]. Thus, $d$ is usually an inaccessibly large quantity, comparable to $D$.
@@
Proposition 3.1 (Quantum eigenstate thermalization implies thermalization o.a. given a nondegenerate spectrum [10, 15]).
@@
Though well known [10, 15, 17, 18, 19], we briefly review the proof in App. D.2.1, as we will use similar arguments on other occasions in this paper. It is based on the following relation: If the energy levels $E_{n}$ within $\Sigma_{d}$ are nondegenerate, then for any observable $\hat{\Pi}_{A}:\mathcal{H}\to\mathcal{H}$ in the full system, and any density operator $\hat{\rho}:\Sigma_{d}\to\Sigma_{d}$ in the energy shell,
@@ 3.2 Eigenspace thermalization $\implies$ idealized thermalization
Figure 4: Schematic depiction of eigenspace thermalization in terms of the relevant matrix elements between pairs $(n,m)$ of the energy levels $(E_{n},E_{m})$; essentially, only the $E_{n}=E_{m}$ pairs are relevant. This matrix element structure can appear different between a spectrum with rare/small degeneracies (4(a)) and with frequent/large degeneracies (4(b)), and therefore more complicated than eigenstate thermalization. However, in terms of energy differences it corresponds to the extremely natural $\Delta E=0$ independent of the spectral details, and can imply thermalization without any knowledge of the spectrum. Even in this case, the set of pairs with $\Delta E=0$ is still not accessible, and as we will argue in Sec. 4.2, eigenspace thermalization is not necessary for accessible thermalization especially in highly degenerate systems.
@@
for some small $\epsilon$. Such a relation has been used to define thermalization in eigenspaces in, e.g., Ref. [50].
While the $\lvert\psi\rangle$ are dynamically independent, they are not all linearly independent; it is therefore desirable to re-express Eq. (45) in terms of a simpler, more tractable criterion in $\mathcal{H}(E)$. For this purpose, we note that the Cauchy-Schwarz inequality (applied to the trace inner product $\left\lvert\operatorname{Tr}[\hat{P}^{\dagger}\hat{Q}]\right\rvert\leq\sqrt{\operatorname{Tr}[\hat{P}^{\dagger}\hat{P}]\operatorname{Tr}[\hat{Q}^{\dagger}\hat{Q}]}$ for $\hat{P}=\lvert\psi\rangle\langle\psi\rvert$ and $\hat{Q}=\hat{\Pi}(E)\left(\hat{\Pi}_{A}-\langle\hat{\Pi}_{A}\rangle_{\Sigma_{d}}\hat{\mathds{1}}\right)\hat{\Pi}(E)$) implies:
@@ 4.1 Weighted time averages
We will also take $w(t)\geq 0$. Among other things, through a relation between non-negativity conditions and Fourier transforms [71, 72], this implies that
@@
Finally, we need to formalize the notion of when $w(t)$ represents a “long-time average”, relative to the timescale set by the bandwidth $\Delta E$ For this, we want $w(t)$ to have appreciable magnitude at least over some long time interval $T\sim 2\pi/\Delta E$, such that $\widetilde{w}(\delta E)$ is appreciable only within $\lvert\delta E\rvert<\Delta E$. Specifically, for some small cutoff $0<w_{0}<1$, we require that $\widetilde{w}(\delta E)$ satisfies:
@@ 4.2 Thermalization in energy bands
$\hat{A}]_{\Sigma_{d},\Delta E}\equiv\sum_{\begin{subarray}{c}n,m:\\ \lvert E_{n}\rangle,\lvert E_{m}\rangle\in\Sigma_{d}\\ \lvert E_{n}-E_{m}\rvert<\Delta E\end{subarray}}\lvert E_{n}\rangle\langle E_{n}\rvert\hat{A}\lvert E_{m}\rangle\langle E_{m}\rvert$. (61)
@@
We can now define “energy-band thermalization” [^12] We note that this structure also occurs if one combines semiclassical results [40, 41] on diagonal and off-diagonal matrix elements connected to classically ergodic systems, and formally removes the $\Delta E\to 0$ limit. by analogy with eigenspace thermalization [Definition 3.2], with the pair of eigenspace projectors $\hat{\Pi}(E)$ replaced by the energy-band selector in Eq. (60):
@@
However, for this to be nontrivial with “accessible” values of $\epsilon$, the number of distinct energy levels in the energy shell should be close to its dimension, $n\gg\epsilon^{2}d/\lambda^{2}$, so that $n_{\lambda}\ll n$ by the above constraint. In that case, we can claim “weak eigenspace thermalization” ($n_{\lambda}/n\ll 1$), in analogy with the notion of “weak ETH” [34, 35, 18] where almost all (but not necessarily all) eigenstates thermalize; more discussion on the physical relevance of such “weak” notions will follow in the Conclusion, Sec. 9.
However, there is a complication: the number of eigenspaces is guaranteed to be accessible in an experiment only if $n$ itself is accessibly small [^13] This is because in a system with $n$ eigenspaces, the spectral form factor [62] (see also Sec. 8), used to measure energy level statistics [22, 23], may oscillate at late times around values as small as $1/n$., such as $n<cN^{\alpha}$ for an $N$-particle system, which requires an unusually high level of degeneracy (close to $\Theta(d)$) in most energy levels. It is therefore difficult to establish that a system has the requisite high number of eigenspaces in most cases.
@@
    If one wants to keep to “accessible” values of $\epsilon$, one can only show weak eigenspace thermalization, that too in a (formally) restricted class of systems where the number of levels still scales as nearly $d$ (i.e. where a high degree of degeneracy is not typical), though this may cover several cases of interest (e.g., such as systems with a time reversal symmetry showing Kramers’ twofold degeneracy [62], but otherwise assumed to be nondegenerate).
@@ 4.3 Energy-band thermalization $\implies$ almost all physical states thermalize
for energy-band thermalization to guarantee thermalization with accuracy $\lambda$, provided that $\mu(\rho)$ is “accessible”. This is in fact anticipated by classical mechanics: such states are precisely the quantum analogue of the “bulky” states in Sec. 2. This analogy is quite strong [^14] We have sometimes been tempted to refer to this (old and well established) semiclassical relation as “Entanglement = Phase space volume”, with the qualifier that entanglement here means the exponential of the second Rényi entropy, following a set of conjectures in quantum gravity [73] relating quantum information measures to more geometric (semi-)classical ones.: if a classical limit exists, the parameter $\mu(\rho)$ can be directly identified with the phase space volume of a classical region in which (say) $\rho$ has uniform support (if it is associated with such a classical state) [68, 69].
@@
$\mathcal{B}=\{\lvert k\rangle\}_{k=0}^{d-1}\;\;\ \subset\Sigma_{d}.$ (71)
@@
$\mathcal{B}=\left\{\lvert s_{1},\ldots,s_{N}\rangle:s_{j}\in\{{0,1}\}\right\},$ (72)
@@
Then, for any complete orthonormal basis of pure states $\mathcal{B}$ within the energy shell $\Sigma_{d}$, the fraction $f_{\lambda}[\mathcal{B}]$ of states that fail to thermalize o.a. to accuracy $\lambda$, i.e. [with $\overline{\Theta}(x)=1$ if $x$ is true and 0 otherwise]
@@
As long as $\epsilon,w_{0}\ll\lambda$, the fraction of basis states with non-thermal time averages to accuracy $\lambda$ (weighted by a given $w(t)$) is vanishingly small. But we cannot force $f_{\lambda}[\mathcal{B}]<1/d$ (which would ensure that all states thermalize) with accessible values of $\epsilon$ and $w_{0}$: there appears to be no accessible way to rule out a small fraction of non-thermal physical states in any accessible time average. These are a quantum analogue of measure-zero exceptions, such as periodic orbits in a classical system showing measure-theoretic ergodicity. As noted in Eq. (68), it is possible to rule out these features with “inaccessible” values of $\epsilon$ and $w_{0}$ by directly considering the thermalization of pure states given energy-band thermalization, provided that such small values of $\epsilon$ are allowed for $\hat{\Pi}_{A}$ by the system.
@@ 5 Global thermalization from a single initial state
To connect autocorrelators to global energy-band thermalization, we only need a straightforward adaptation of the proofs of wavefunction ergodicity [40, 41, 46] from global eigenstate thermalization to energy-band thermalization with a careful accounting of weighted time averages. This is carried out in Sec. 5.2. More significantly, in Sec. 5.3, we combine this result with those of Sec. 4 to completely bypass properties of the energy levels, and directly connect the decay of autocorrelators to global thermalization, realizing a fully quantum analogue of the classical statement containing Eq. (1).